1. Write the best Lewis structure (minimize formal charge, maximize number of resona structures) for the following species.
Answer #1

Ans: Lewis Structure- NOTA - Ñ =Ö Es Ö=ÑOLA FC = (Valence Shell electrons of Atom) - (Numbes on bonds Associated with the AtoElectronic geometry - linear :C=0: linear Molecular geometry - polar Molecule So - 10 :0:0 :00 F.C = (0,2 = 6-1-6=- 1 - (01)Poyo to-=0 ( 20--0 <>03BÖ: 0 :00 FC = 6-1-6--1. (oxygen 2)) F. C = 6-1-6= -1 Coxygen 2 FC:6-1-6--1 (oxygen ) Electronic geomeg. Better Lewis Structure - 1 - Ć - S - H formal charge for better (Minimize the Lewis Structure)